# Wrinkling and imaging of thin curved sheets

### v1

@@ V Results and Discussion
Figure 8: Reconstructed surface maps for floating curved sheets of various inherent Gaussian curvatures $\kappa=1/R_{1}R_{2}$. The top row shows films cut from spherical surfaces with radii of $R_{1}=R_{2}=$ 30, 40, 50, 60 and 70 mm. Films with negative Gaussian curvature were cut from saddle-shaped surfaces with different $R_{1}$, and fixed $R_{2}=-70$ mm. The middle row shows the wrinkle patterns for $R_{1}=$ 40, 50, 60, 70, and 80 mm, while the bottom row shows the patterns for $R_{1}=$ 90, 100, 110, 120 and 130 mm.
@@
Figure 10: Wrinkle amplitudes are related to a film’s natural Gaussian curvature, and to the excess of its surface area as compared to a flat cross-sectional profile. (a) The relative excess area as estimated from the wrinkled film’s RMS amplitude, $2\pi^{2}A^{2}/\lambda^{2}$, can be compared to the excess area of the undeformed shell, $|\kappa|W^{2}/12$. While these measures are proportional to each other, the sign of $\kappa$ has an effect on how the area is distributed: the spherical shells (blue) show larger amplitudes and areas, compared to the hyperboloid shells (red). (b) For the negatively curved shells, systematically measuring the total surface area, $\mathcal{A}$, of the wrinkled and undeformed configurations shows that there is a deficit in relative excess area, $\mathcal{A}/W^{2}-1$, for the wrinkled films. (c) For the negatively curved shells, the wrinkle amplitude increases with distance away from the center of the film, and with $\sqrt{\kappa}$. There are, however, flat regions of width $\sim 5$ mm around the film center, and the stable lines.
Figure 8 gives the reconstructed surface profiles of all the films used in this study. As can be seen for the full set of experiments, our representative descriptions of the wrinkled domains hold for shells with a wide range of Gaussian curvatures. In general, we find that varying the magnitude of $\kappa$ only changes the amplitude of the resulting pattern, without otherwise affecting its structure. The films with positive curvature, $\kappa>0$, all show a disordered central region, with a few diagonal wrinkles appearing in the corners of the film. The films with negative curvature, $\kappa<0$, all show flat regions along the stable lines, separating four distinct domains of well-ordered wrinkles. These consistent features allow us to develop robust measures of wrinkle wavelength and amplitude, extracted from the surface profile data.
These results also demonstrate that the wrinkles depend only on the Gaussian curvature, rather than on the individual values of the principle curvatures $\kappa_{1}$ and $\kappa_{2}$. For the negatively curved films shown in Fig. 8 these are the curvatures that would be measured along the perpendicular edges of a film, and can differ from each other by up to about a factor of two. There are no corresponding differences in the wavelength, amplitude, or positioning of the wrinkles that appear along these directions, however, and no evidence of symmetry-breaking due to the differences in principle curvatures. Although expected from Gauss’ theorema egregium [22, 44], this remains an unintuitive and remarkable result.
The amplitude of the wrinkle patterns, and the excess surface area that these features accommodate, are explored in Fig. 10. As an initial measure of the amplitude of a wrinkle pattern we use the RMS amplitude of the surface distortion, $A=\sqrt{\langle|h-h_{0}|^{2}\rangle}$, measured over the entire film. If the wrinkles had a uniform sinusoidal cross-section of wavelength $\lambda$ and amplitude $A$, then they would have an excess surface area of $2\pi^{2}A^{2}/\lambda^{2}$, relative to a flat film with the same cross-section. The undeformed shell will also have a relative excess of area, given by $|\kappa|W^{2}/12$. These measures both represent a type of misfit strain between the natural surface area of the film, and the cross-sectional area of the fluid bath that it covers. They are derived from Eqs. 8 and 9, and compared to each other in Fig. 10(a). For the positively curved films, where the assumption of a constant amplitude is more plausible (see top row of Fig. 8) a least-squares regression gives a slope of 0.95$\pm$0.21. This suggests that the film’s surface area is reasonably well conserved as it wrinkles, in line with the assumptions of most models [22, 29, 24]. For the negatively curved shells, however, the best-fit line between these two measures of excess area has a slope of only $0.27\pm 0.04$. This is a clear difference in how the sign of the curvature affects the wrinkle amplitude.
For the negatively curved shells, the wrinkling amplitude varies considerably over the film, with flat areas surrounding the stable lines. To more carefully evaluate these shells, we used the surfarea function [46] in Matlab to measure the total surface area of the reconstructed height profiles of their wrinkled surfaces. From these areas $\mathcal{A}$ we directly calculated an excess area $\mathcal{A}/W^{2}-1$, relative to the area of the water which they cover, $W^{2}$. We used the same methods to measure the relative excess area of hyperboloid surfaces with constant curvatures, for comparison, with the same lateral dimensions as our molds. As shown in Fig. 10(b), the wrinkled films made from negatively curved surfaces have considerably less surface area than the reference, undeformed shells. Specifically, while the reference shells show good agreement with the expected limit of $\mathcal{A}/W^{2}-1=|\kappa|W^{2}/12$, these two measures of the relative excess area differ by a best-fit factor of $0.50\pm 0.04$ for the wrinkled configurations. In other words, these floating curved sheets have lost about half of their excess area.
Additional evidence of some film compression, for the negatively-curved shells, can be seen in Fig. 10(c). This plot shows how the RMS amplitude varies from the center of these films, to their edges edges. The averages are calculated on concentric square boundaries, along lines a fixed distance from the film’s center. Over much of the film the wrinkle amplitude increases smoothly with the distance from the center (practically, the distance from a stable line). However, there is a flattened region of width $\sim 5$ mm near the center of the film, or around the stable lines, which is not predicted by the current models [29, 24]. In these regions, with vanishingly small wrinkle amplitude, local film compression would appear to be required.
@@ VI Summary and Conclusions
To achieve these results, we developed an imaging technique based on the methods of synthetic Schlieren imaging, which have previously proved useful for generating topographical maps of the height $h$ of deformed liquid surfaces [31, 30, 15, 33]. Our approach does not require any assumption of small surface perturbations, i.e. it avoids the constraint that $(h-h_{0})\ll h_{0}$, where $h_{0}$ is the reference height of the imaged surface. The change of variables involved in this approach also circumvents the issue of over-constrained data encountered when using two scalar fields, namely the displacement of visual features in two independent directions, to reconstruct the single scalar field of the surface height. We further demonstrated how the use of a series of projected features can increase the accuracy of the reconstructed surface, without introducing ambiguities in feature matching, and provided estimates of this accuracy, based on the density of sampling points and the Nyquist sampling condition.
Together, these results provide insight into the effects of curvature on wrinkling phenomena, and an imaging method with broad applicability. Since the stiffness introduced by a buoyant substrate is simply a linear restoring force, these findings are also appropriate to rigid films attached to soft elastic substrates, where there is some mismatch in curvature between the film and substrate. This could provide for applications in shape memory, for example. Other relevant curved films and interfaces are encountered in diverse settings, including in the brain [20] and gut [26], swelling hydrogels [19], liquid lenses [33], deformable [47] or reconfigurable [48] microfluidics, stretchable [49] and reconfigurable [50] electronics, flexible optical components such as solar panels and diffraction gratings [51], and graphene sheets [52].

### v2

@@ V Results and Discussion
Figure 8: Reconstructed surface maps for floating curved sheets of various inherent Gaussian curvatures $\kappa=1/R_{1}R_{2}$. The top row shows films cut from spherical surfaces with radii of $R_{1}=R_{2}=$ 30, 40, 50, 60 and 70 mm. Films with negative Gaussian curvature were cut from saddle-shaped surfaces with different $R_{1}$, and fixed $R_{2}=-70$ mm. The middle row shows the wrinkle patterns for $R_{1}=$ 40, 50, 60, 70, and 80 mm, while the bottom row shows the patterns for $R_{1}=$ 90, 100, 110, 120 and 130 mm. In the right, upper two panels, dashed lines indicate the locations of the cross-sectional profiles shown in Fig. 9
Figure 9: Cross-sectional profiles of floating shells with (a) positive and (b) negative intrinsic curvature. Profiles are extracted at the locations shown by the dashed lines in Fig. 8. Red and blue shading highlight locations that belong to different domains, as predicted by the method of stable lines (see Fig. 7).
@@
Figure 8 gives the reconstructed surface profiles of all the films used in this study, with some cross-sectional profiles across films also given in Fig. 9. As can be seen for the full set of experiments, our representative descriptions of the wrinkled domains hold for shells with a wide range of Gaussian curvatures. In general, we find that varying the magnitude of $\kappa$ only changes the amplitude of the resulting pattern, without otherwise affecting its structure. The films with positive curvature, $\kappa>0$, all show a disordered central region, with a few diagonal wrinkles appearing in the corners of the film. The films with negative curvature, $\kappa<0$, all show flat regions along the stable lines, separating four distinct domains of well-ordered wrinkles. These consistent features allow us to develop robust measures of wrinkle wavelength and amplitude, extracted from the surface profile data.
These results also demonstrate that the wrinkles depend only on the Gaussian curvature, rather than on the individual values of the principal curvatures $\kappa_{1}$ and $\kappa_{2}$. For the negatively curved films shown in the lower two rows of Fig. 8, these are the curvatures that would be measured along the perpendicular edges of a film, and can differ from each other by up to about a factor of two. There are no corresponding differences in the wavelength, amplitude, or positioning of the wrinkles that appear along these directions, however, and no evidence of symmetry-breaking due to the differences in principal curvatures. Although expected from Gauss’ theorema egregium [22, 44], this remains an unintuitive and remarkable result.
The amplitude of the wrinkle patterns, and the excess surface area that these features accommodate, are explored in Fig. 11. As an initial measure of the amplitude of a pattern we use the RMS amplitude of the surface distortion, $A=\sqrt{\langle|h-h_{0}|^{2}\rangle}$, measured over the entire film. If the wrinkles had a uniform sinusoidal cross-section of wavelength $\lambda$ and amplitude $A$, then they would have an excess surface area of $2\pi^{2}A^{2}/\lambda^{2}$, relative to a flat film with the same cross-section. The undeformed shell will also have a relative excess of area, given by $|\kappa|W^{2}/12$. These measures both represent a type of misfit strain between the natural surface area of the film, and the cross-sectional area of the fluid bath that it covers. They are derived from Eqs. 8 and 9, and compared to each other in Fig. 11(a). For the positively curved films, where the assumption of a constant amplitude is more plausible (see top row of Fig. 8) a least-squares regression gives a slope of 0.95$\pm$0.21. This suggests that the film’s surface area is reasonably well conserved as it wrinkles, in line with the assumptions of most models [22, 29, 24]. For the negatively curved shells, however, the best-fit line between these two measures of excess area has a slope of only $0.27\pm 0.04$. This is a clear difference in how the sign of the curvature affects the wrinkle amplitude.
For the negatively curved shells, the wrinkling amplitude varies considerably over the film, with flat areas surrounding the stable lines. To more carefully evaluate these shells, we used the surfarea function [46] in Matlab to measure the total surface area of the reconstructed height maps of their wrinkled surfaces. From these areas $\mathcal{A}$ we directly calculated an excess area, $\mathcal{A}/W^{2}-1$, relative to the area of the water which they cover, $W^{2}$. We used the same methods to measure the relative excess area of hyperboloid surfaces with constant curvatures, for comparison, with the same lateral dimensions as our molds. As shown in Fig. 11(b), the wrinkled films made from negatively curved surfaces have considerably less surface area than the reference, undeformed shells. Specifically, while the reference shells show good agreement with the expected limit of $\mathcal{A}/W^{2}-1=|\kappa|W^{2}/12$, these two measures of the relative excess area differ by a best-fit factor of $0.50\pm 0.04$ for the wrinkled configurations. In other words, these floating curved sheets have lost about half of their excess area.
Additional evidence of some film compression, for the negatively-curved shells, can be seen in Fig. 11(c). This plot shows how the RMS amplitude varies from the center of these films, to their edges. The averages are calculated on concentric square boundaries, along lines a fixed distance from the film’s center. Over much of the film the wrinkle amplitude increases smoothly with the distance from the center (practically, the distance from a stable line). However, there is a flattened region of width $\sim 5$ mm near the center of the film, or around the stable lines, which is not predicted by the current models [29, 24]. In these regions, with vanishingly small wrinkle amplitude, local film compression would appear to be required.
@@
Figure 11: Wrinkle amplitudes are related to a film’s natural Gaussian curvature, and to the excess of its surface area as compared to a flat cross-sectional profile. (a) The relative excess area, as estimated from a wrinkled film’s RMS amplitude, $2\pi^{2}A^{2}/\lambda^{2}$, can be compared to the excess area of the undeformed shell, $|\kappa|W^{2}/12$. While these measures are proportional to each other, the sign of $\kappa$ has an effect on how the area is distributed: the spherical shells (blue) show larger amplitudes and areas, compared to the hyperboloid shells (red). (b) For the negatively curved shells, systematically measuring the total surface area, $\mathcal{A}$, of the wrinkled and undeformed configurations shows that there is a deficit in relative excess area, $\mathcal{A}/W^{2}-1$, for the wrinkled films. (c) For the negatively curved shells, the wrinkle amplitude increases with distance away from the center of the film, and with $\sqrt{\kappa}$. There are, however, flat regions of width $\sim 5$ mm around the film center, and the stable lines.
@@ VI Summary and Conclusions
To achieve these results, we developed an imaging technique based on the methods of synthetic schlieren imaging, which have previously proved useful for generating topographical maps of the height $h$ of deformed liquid surfaces [31, 30], including for cases of a floating annulus [15] or drying droplets [33]. This approach relies on measurements of the refraction of rays as they pass through a deformed, wrinkled surface.
Our approach does not require any assumption of small surface perturbations, i.e. it avoids the constraint that $(h-h_{0})\ll h_{0}$, where $h_{0}$ is the reference height of the imaged surface. The change of variables involved in this approach also circumvents the issue of over-constrained data encountered when using two scalar fields, namely the displacement of visual features in two independent directions, to reconstruct the single scalar field of the surface height. We further demonstrated how the use of a series of projected features can increase the accuracy of the reconstructed surface, without introducing ambiguities in feature matching, and provided estimates of this accuracy, based on the density of sampling points and the Nyquist sampling condition.
Together, these results provide insight into the effects of curvature on wrinkling phenomena, and an imaging method with broad applicability. Since the stiffness introduced by a buoyant substrate is simply a linear restoring force, these findings are also appropriate to rigid films attached to soft elastic substrates, where there is some mismatch in curvature between the film and substrate; This could provide for applications in shape memory, for example. As PDMS is photoelastic [47], like many other polymers, crossed-polarizers could be used in parallel with these methods, to evaluate local strains and stretching. Local stretching could also be studied through detailed analysis of the perimeter to area ratio of circles inscribed within the film, either physically or by tracing the appropriate paths within the surface of a reconstructed film. Other relevant curved films and interfaces are encountered in diverse settings, including in the brain [20] and gut [26], swelling hydrogels [19], liquid lenses [33], deformable [48] or reconfigurable [49] microfluidics, stretchable [50] and reconfigurable [51] electronics, flexible optical components such as solar panels and diffraction gratings [52], and graphene sheets [53].